OpenStudy (anonymous):

identify the solutions of the system of equations, if any. 2x + 5y = 10 y = 2 - x

OpenStudy (wolf1728):

A) 2x + 5y = 10 y = 2 - x Let's change this equation a little: B) x + y = 2 Multiply B by -2 B) -2x - 2y = -4 then adding this to A A) 2x + 5y = 10 we get 3 y = 6 y = 2 Therefore x = 0